OpenStudy (anonymous):

Water flows with 1 gal/min and 10oF temperature difference. Calculate the ton of cooling load ?

OpenStudy (fretje):

So you need to cool water 10 Fahrenheit grades at 1 gallon per minute. From the web I get http://www.engineeringtoolbox.com/cooling-loads-d_665.html 1 refrigeration ton = 12661 kJ/h. 1000J/kJ.1hour/3600 s = 3516,94 J/s and a (US) gallon is 3,78541E-3 m³ and the heat capacity of water = 4,18 J/kgK.1000kg/1m³ = 4180J/m³K and for dT -> 1K = 1,8°F (for temperature differences: one degree kelvin equals 1,8 degree Fahrenheit) so i calculate x which is the cooling load: x = 1gal/min. 3,78541E-3m³/ 1 gal .1min/60s. 4180J/m³K.1refr ton/3516,94 J/s.10°F.10K/18°F = 4,165368E-4 refr.tons (scientific notation) or thus 0,0004165 refr tons.